OpenStudy (anonymous):

find the product write your answer in exponential form 7^-7*7^-3

OpenStudy (anonymous):

\[7^{-7}*7^{-3}\]

zepdrix (zepdrix):

Rule of exponents tells us to `add` the exponent values, yes? \[\Large\rm 7^{-7}\cdot 7^{-3}=7^{-7+(-3)}\]

OpenStudy (anonymous):

\[8=8^{4}*blank\]

OpenStudy (kirbykirby):

\(8 \) is the same as \(8^1\) so, \(8^1=8^4*x\). We can think of having something like \(8^1 = 8^4 * 8^{\text{blank}}\) Think of which number, when added to 4, will 1

OpenStudy (anonymous):

I need help with more

OpenStudy (kirbykirby):

will give 1*

OpenStudy (anonymous):

\[9^{2}*9^{-6}\]

OpenStudy (kirbykirby):

add the exponents together

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[4^{4}*blank4^{^{4}}\]